derive a reduction formula e^(mx)/x^ndx
1
Expert's answer
2021-06-14T16:20:45-0400

u=emx

du=memxdx

dv=x-ndx

v=x1-n/(1-n)

"\int \frac{e^mx}{x^n}dx=\frac{e^{mx}x^{1-n}}{1-n}-m\int \frac{x^{1-n}e^{mx}}{1-n}dx"